# Black Hole Ringdown Amplitudescopy

### v1

@@ Acknowledgements.
We thank Juan Calderon Bustillo, Gregorio Carullo, Giovanni D’Addario, Matteo Della Rocca, and Jacopo Lestingi for interesting discussions. P.P. and F.C. are partially supported by the MUR FIS2 Advanced Grant ET-NOW (CUP: B53C25001080001) and by the INFN TEONGRAV initiative. F.C. acknowledges the financial support provided under the “Progetti per Avvio alla Ricerca Tipo 1”, protocol number AR12419073C0A82B. X.J. is supported by the Spanish Ministerio de Ciencia, Innovación y Universidades (Beatriz Galindo, BG22-00034) and cofinanced by UIB; the Spanish Agencia Estatal de Investigación Grants No. PID2022-138626NB-I00, No. RED2022- 134204-E, and No. RED2022-134411-T, funded by MCIN/AEI/10.13039/501100011033/FEDER, UE; the MCIN with funding from the European Union NextGenerationEU/PRTR (No. PRTR-C17.I1); the Comunitat Autonòma de les Illes Balears through the Direcció General de Recerca, Innovació I Transformació Digital with funds from the Tourist Stay Tax Law (No. PDR2020/11 - ITS2017-006), and the Conselleria d’Economia, Hisenda i Innovació Grant No. SINCO2022/6719. Some numerical computations have been performed at the Vera cluster supported by the Italian Ministry of Research and by Sapienza University of Rome, with the University of Birmingham’s BlueBEAR and the computer resources at MareNostrum5 (RES-AECT-2025-2-0038).

### v2

@@ Acknowledgements.
We thank Juan Calderon Bustillo, Gregorio Carullo, Giovanni D’Addario, Matteo Della Rocca, Cecilia Maria Fabbri, Stephen Green, Jacopo Lestingi, Laura Sberna, and Thomas Sotiriou for interesting discussions. P.P. and F.C. are partially supported by the MUR FIS2 Advanced Grant ET-NOW (CUP: B53C25001080001) and by the INFN TEONGRAV initiative. F.C. acknowledges the financial support provided under the “Progetti per Avvio alla Ricerca Tipo 1”, protocol number AR12419073C0A82B. X.J. is supported by the Spanish Ministerio de Ciencia, Innovación y Universidades (Beatriz Galindo, BG22-00034) and cofinanced by UIB; the Spanish Agencia Estatal de Investigación Grants No. PID2022-138626NB-I00, No. RED2022- 134204-E, and No. RED2022-134411-T, funded by MCIN/AEI/10.13039/501100011033/FEDER, UE; the MCIN with funding from the European Union NextGenerationEU/PRTR (No. PRTR-C17.I1); the Comunitat Autonòma de les Illes Balears through the Direcció General de Recerca, Innovació I Transformació Digital with funds from the Tourist Stay Tax Law (No. PDR2020/11 - ITS2017-006), and the Conselleria d’Economia, Hisenda i Innovació Grant No. SINCO2022/6719. Some numerical computations have been performed at the Vera cluster supported by the Italian Ministry of Research and by Sapienza University of Rome, with the University of Birmingham’s BlueBEAR and the computer resources at MareNostrum5 (RES-AECT-2025-2-0038).
Appendix A Frequency shift computations
In this appendix, we provide the computation of the frequency shifts described in Refs. [40, 37]. We write the beyond-GR correction to the complex frequency in geometric units as
$\omega_{i}(\chi)=\omega_{i,0}(\chi)+\zeta\,\omega_{i,1}(\chi),$ (13)
where $\omega_{i,0}$ is the GR QNM complex frequency, and $\omega_{i,1}$ is the first-order correction, which we decompose into real and imaginary parts,
$\omega_{i,0}=\Re(\omega_{i,0})+i\,\Im(\omega_{i,0}),\qquad\omega_{i,1}=\Re(\omega_{i,1})+i\,\Im(\omega_{i,1}).$ (14)
We define the spin-dependent functions of Eq. (7) as
$F_{i}(\chi)\equiv\frac{\Re(\omega_{i,1}(\chi))}{\Re(\omega_{i,0}(\chi))},\qquad T_{i}(\chi)\equiv\frac{\Im(\omega_{i,1}(\chi))}{\Im(\omega_{i,0}(\chi))},$ (15)
so that
$\Re(\omega_{i,1})=F_{i}(\chi)\,\Re(\omega_{i,0}),\qquad\Im(\omega_{i,1})=T_{i}(\chi)\,\Im(\omega_{i,0}).$ (16)
It follows that the corrected real and imaginary parts can be written as
$\Re(\omega_{i})$ $=\Re(\omega_{i,0})\bigl[1+\zeta\,F_{i}(\chi)\bigr],$ (17)
$\Im(\omega_{i})$ $=\Im(\omega_{i,0})\bigl[1+\zeta\,T_{i}(\chi)\bigr].$ (18)
The physical frequency and damping time are defined by
$f_{i}\equiv\frac{\Re(\omega_{i})}{2\pi},\qquad\tau_{i}\equiv-\frac{1}{\Im(\omega_{i})},$ (19)
with the GR values
$f_{i,0}=\frac{\Re(\omega_{i,0})}{2\pi},\qquad\tau_{i,0}=-\frac{1}{\Im(\omega_{i,0})}.$ (20)
Therefore,
$f_{i}$ $=f_{i,0}\bigl[1+\zeta\,F_{i}(\chi)\bigr],$ (21)
$\tau_{i}$ $=\tau_{i,0}\,\frac{1}{1+\zeta\,T_{i}(\chi)}.$ (22)
Defining the relative deviations
$\delta f_{i}\equiv\frac{f_{i}}{f_{i,0}}-1,\qquad\delta\tau_{i}\equiv\frac{\tau_{i}}{\tau_{i,0}}-1,$ (23)
we obtain the exact relations
$\delta f_{i}$ $=\zeta\,F_{i}(\chi),$ (24)
$\delta\tau_{i}$ $=-\,\frac{\zeta\,T_{i}(\chi)}{1+\zeta\,T_{i}(\chi)}$ (25)
which are the terms of Eq. (7). The first-order complex correction $\omega_{i,1}(\chi)$ is expressed as a polynomial in spin:
$\omega_{i,1}(\chi)=\sum_{j=0}^{8}\omega^{(i)}_{j}\,\chi^{j},\qquad\omega^{(i)}_{j}=a^{(i)}_{j}+i\,b^{(i)}_{j}$, (26)
where the complex coefficients $\omega^{(i)}_{j}$ are obtained from the fits in Table V of Ref. [40] for GB, and Table V of Ref. [37] for CS. The tables below report the full complex coefficients $\omega^{(i)}_{j}$ for the dominant $(2,2,0)$ and $(3,3,0)$ modes in CS and GB, separately for the axial and polar parities.
220220 330330
$j$ axial polar axial polar
0 0.2460410.125482ii-0.246041-0.125482i\,i 0 0.9127520.164615i-0.912752-0.164615\,i 0
1 0.359080.184144i-0.35908-0.184144\,i 0.0307839+0.017411i-0.0307839+0.017411\,i 1.095120.0911042i-1.09512-0.0911042\,i 0.09015350.000837525i-0.0901535-0.000837525\,i
2 0.595672+0.0876598i-0.595672+0.0876598\,i 0.0138350.214353i0.013835-0.214353\,i 1.23272+0.116646i-1.23272+0.116646\,i 0.07635260.149877i-0.0763526-0.149877\,i
3 1.41373+0.258192i1.41373+0.258192\,i 0.3987610.4473760.398761-0.447376 2.37515+0.763688i2.37515+0.763688\,i 0.4520640.248466i0.452064-0.248466\,i
4 11.2668+0.703524i-11.2668+0.703524\,i 6.66087+0.448166i-6.66087+0.448166\,i 16.74270.927771i-16.7427-0.927771\,i 6.706640.508568i-6.70664-0.508568\,i
5 42.0081.28144i42.008-1.28144\,i 24.73621.091i24.7362-1.091\,i 52.5812+1.12866i52.5812+1.12866\,i 25.7745+1.18438i25.7745+1.18438\,i
6 82.8112+1.05377i-82.8112+1.05377\,i 55.9612+2.5070i-55.9612+2.5070\,i 97.6497+0.865323i-97.6497+0.865323\,i 60.52451.60982i-60.5245-1.60982\,i
7 83.75080.833429i83.7508-0.833429\,i 63.1241.10815i63.124-1.10815\,i 97.29531.91376i97.2953-1.91376\,i 60.52451.60982i-60.5245-1.60982\,i
8 34.8784-34.8784 28.7982-28.7982 40.6093-40.6093 32.9995-32.9995
Table 2: Complex coefficients $\omega^{(i)}_{j}$ for the $(220)$ and $(330)$ modes and axial/polar parities in CS theory.
220220 330330
$j$ axial polar axial polar
0 0.055241+0.00686713i0.055241+0.00686713\,i 0.2152020.0734094i-0.215202-0.0734094\,i 0.109706+0.0061152i0.109706+0.0061152\,i 0.8727890.113506i-0.872789-0.113506\,i
1 0.985294+0.0636233i0.985294+0.0636233\,i 2.518160.411031i-2.51816-0.411031\,i 0.685561+0.231546i0.685561+0.231546\,i 1.201980.339974i-1.20198-0.339974\,i
2 18.4902+0.325569i-18.4902+0.325569\,i 48.4695+9.59898i48.4695+9.59898\,i 11.67334.20702i-11.6733-4.20702\,i 2.82484+9.09087i2.82484+9.09087\,i
3 157.8024.46276i157.802-4.46276\,i 431.42882.4765i-431.428-82.4765\,i 108.698+37.6761i108.698+37.6761\,i 35.602475.7708i-35.6024-75.7708\,i
4 682.224+19.8274i-682.224+19.8274\,i 1935.01+378.832i1935.01+378.832\,i 510.61178.723i-510.61-178.723\,i 165.435+341.913i165.435+341.913\,i
5 1660.0855.4163i1660.08-55.4163\,i 4831.15965.058i-4831.15-965.058\,i 1334.38+467.415i1334.38+467.415\,i 434.981857.681i-434.981-857.681\,i
6 2307.78+87.744i-2307.78+87.744\,i 6808.99+1386.33i6808.99+1386.33\,i 1970.88687.899i-1970.88-687.899\,i 638.737+1219.16i638.737+1219.16\,i
7 1711.3273.0696i1711.32-73.0696\,i 5065.11050.91i-5065.1-1050.91\,i 1536.55+532.083i1536.55+532.083\,i 488.58915.499i-488.58-915.499\,i
8 526.516+25.0324i-526.516+25.0324\,i 1544.32+325.905i1544.32+325.905\,i 493.701167.701i-493.701-167.701\,i 150.043+280.837i150.043+280.837\,i
Table 3: Same of Table 2, but for GB theory.
Appendix B Additional results
In this subsection, we report additional results of the analyses performed. In Fig. 4, we report the analog of the top-left panel of Fig. 2 (i.e. the CS case), but including $\gamma$ in the sampling parameters. In Table 4 instead, we report the analog of Table 1 for the CS case, where $\gamma$ is sampled. Further, in Figs. 5-8 we report the full corner plots related to Fig. 3, for the cases of $\gamma_{\rm inj}=0$ and $\gamma_{\rm inj}=10$.
Figure 4: Same as top left panel of Fig. 2, but sampling on $\gamma$.
CS 0.06 -0.42 -1.75
Table 4: Same as first row of Tab. 1, but for the analysis where $\gamma$ is sampled.
Figure 5: Corner plot for GB 220P+330P injection at ${\rm SNR}=80$ with $\sqrt{\alpha}=12$ and $\gamma_{\rm inj}=0$. Curves show posteriors for three injected amplitudes $A_{330}\in\{0.005,\,0.0351,\,0.4\}$ (blue/red/green). Solid curves correspond to recovery with $\gamma=\gamma_{\rm inj}$, dashed curves to recovery with $\gamma=0$. Black dashed lines indicate injected values. For the 2D distributions, we report the 90% credible level.
Figure 6: Same as Fig. 5, but for $\gamma_{\rm inj}=10$.
Figure 7: Same as Fig. 5, but for CS and $\gamma_{\rm inj}=0$.
Figure 8: Same as Fig. 5, but for CS and $\gamma_{\rm inj}=10$.